OpenStudy (anonymous):

whats the order of this function pole at z=-1 h(z)=3/(1-cos((z+1)^2)) and g(z)= (z+1)/)(z^3+1)^4)

OpenStudy (anonymous):

Hint, but not a proof \[ h(z) = \frac{3}{(z+1)^2}+1+\frac{1}{5} (z+1)^2+\frac{2}{63} (z+1)^4+O\left((z+1)^5\right)\\ g(z)=\frac{1}{81 (z+1)^3}+\frac{4}{81 (z+1)^2}+\\ \frac{26}{243 (z+1)}+\frac{40}{243}+\frac{145 (z+1)}{729}+\frac{16}{81} (z+1)^2+O\left((z+1)^3\right) \]

OpenStudy (anonymous):

Read this, you might find it helpful http://www.physicsforums.com/showthread.php?t=609593